If one root of the equation x^2-10x+16=0 is half of the one of the roots of x^2-4Rx+8=0. Find the value of R if it is known that both the equation have integral roots.? Someone please help
Answer by rothauserc(4718)

we are given
x^2-10x+16=0
factor equation
(x-2)*(x-8) = 0
roots are 2 and 8
we are given
x^2-4Rx+8=0 and one of the roots of the first equation is 1/2 of one of the root of this equation, therefore the roots are 4, 2 and
6 = 4R
R = 6/4